OpenStudy (anonymous):

will some one plz help In how many different ways can the following positions be chosen from the 18-member math team? president, vice president, and treasurer

OpenStudy (anonymous):

x^2-2x=-1, 3x-x-x+a-a=5, (x^2-1)/(x+1), 2x-(x+x)

OpenStudy (anonymous):

amm so how do i do that

OpenStudy (anonymous):

You use permutations. In your calculator you would use nPr where n is the total number and r is the positions you're looking to fill. So 16P4 = 43680 The formula for permutations is n!/(n-r)! = n(n-1)(n-2)...(n-r +1) = 16*15*14*13= 43680.

OpenStudy (anonymous):

oo ok ya ya i gt it thanx